# EgoTraj-Bench: Towards Robust Trajectory Prediction Under Ego-view Noisy Observations

### v1

@@ V-B Quantitative Results
We present performance comparisons across all models on both datasets in Table II. In the T2FPV-ETH dataset, our method achieves state-of-the-art performance with minADE@20 of 0.60 and minFDE@20 of 0.74, outperforming the best baseline by over 11% and 15%, respectively. Moreover, all methods exhibit significantly degraded performance compared to their results under clean historical trajectories, confirming that existing BEV-based approaches struggle to handle the realistic perception noise present in ego-centric settings.
On the EgoTraj-TBD dataset, our model shows consistent improvements over existing methods. Notably, as shown in Table III, our model achieves significant improvements when generating fewer future trajectory candidates (i.e., smaller $K$). Specifically, with the same reduced number of samples, our approach improves minADE and minFDE by around 16% compared to the SOTA baseline. The strong performance indicates a predicted distribution more closely aligned with true trajectories, underscoring enhanced robustness and predictive efficiency under noisy conditions.
Additionally, while integrating CoFE brings improvements for some models, the gains are limited. And in flow-based models, performance even degrades. This suggests that correcting only missing positions is insufficient, as ego-centric observations contain diverse noise beyond occlusion, such as tracking errors and perspective distortion. Effective denoising requires holistic trajectory modeling rather than patchwise correction, which is validated by our SOTA results.
Note that for each baseline model, including BiFlow, the ADE/FDE values are lower than those evaluated in T2FPV-ETH. A key reason is our higher-fidelity ground truth and more noise-aware processing pipeline, which results in significantly lower historical MSE (0.66m) in EgoTraj-TBD, as shown in Table I. These more meaningful historical representations enable more reliable training and contribute to overall better performance across methods.
@@ V-C Ablation Study
Table III presents an ablation study of key components in BiFlow: Social Interaction (SI) within the contextual encoder, EgoAnchor (EA) distillation, and Shared Encoder (SE) within the dual stream. We compare against MoFlow as a strong baseline. When only SI is added, BiFlow achieves notable improvements over MoFlow, reducing minADE/minFDE by over 9% at KK=1. Incorporating EA further enhances performance, yielding a 13% improvement with the same $K$. With all components (SI, EA, and SE) integrated, the full model achieves the best performance, improving minADE and minFDE by 16%, 17%, and 16% at KK=1, KK=5, and KK=10, respectively. These results demonstrate that integrating social interaction cues and ego-centric intent patterns, particularly through EgoAnchor, significantly improves prediction accuracy and robustness under the deployment-realistic condition.
Figure 4: Qualitative Results. Solid lines represents the ground truth trajectories, while dashed lines shows the predicted trajectories.
@@ V-D Qualitative Results
Fig. 4 visualizes predicted trajectories on EgoTraj-TBD datasets, demonstarting our BiFlow model produces accurate and physically plausible predictions, particularly under realistic ego-view perturbations such as occlusion, mis-tracked and ego-motion drift.
@@ VI Conclusion
While our approach shows competitive performance in the current setup, its generalization to other platforms may be limited by differences in camera height, field of view, and sensor characteristics, which influence the nature of perception noise. Future work will focus on adapting the framework for diverse robotic embodiments and environmental conditions. We hope that EgoTraj-Bench and BiFlow will support the community with models more robust to the ego-centric observation challenges and progress toward reliable real-world deployment.

### v2

@@ V-B Quantitative Results
We present performance comparisons across all models on both datasets in Table II. Across architectures, ego-view noise substantially degrades performance relative to clean-history results. Recurrent models (VRNN/AC-VRNN/SGNet) degrade more severely due to error accumulation under corrupted histories, while the attention-based predictor TUTR is generally more competitive but still struggles under realistic noise. Among prior methods, the flow-matching model MoFlow is the strongest competitor, suggesting that conditional generative modeling is promising for forecasting from noisy histories, yet it still degrades under corrupted observations.
Under ego-view noise, our method is both more robust and more efficient. On T2FPV-ETH, it achieves state-of-the-art (SOTA) performance with minADE@20 of 0.60 and minFDE@20 of 0.74, outperforming the prior SOTA by over 11% and 15%, respectively. On EgoTraj-TBD, it shows consistent improvements over existing methods. Notably, as shown in Table III, our model achieves significant improvements when generating fewer future trajectory candidates (i.e., smaller $K$). Specifically, with the same reduced number of samples, our approach improves minADE and minFDE by around 16% compared to the prior SOTA. The strong performance indicates a predicted distribution more closely aligned with true trajectories, underscoring enhanced robustness and predictive efficiency under noisy conditions.
Additionally, while integrating CoFE brings improvements for some models, the gains are limited, and flow-based methods can even degrade. This suggests that correcting only missing positions is insufficient, as ego-centric observations contain structured noise beyond occlusion, including tracking errors and perspective distortion. Effective denoising therefore requires holistic trajectory modeling rather than patchwise correction, which is validated by our SOTA results.
Note that for all models, including BiFlow, the ADE/FDE values on EgoTraj-TBD are lower than those evaluated on T2FPV-ETH. A key reason is our higher-fidelity ground truth and more noise-aware processing pipeline, which results in significantly a lower historical MSE (0.66m) in EgoTraj-TBD, as shown in Table I. These more meaningful historical representations enable more reliable training and contribute to overall better performance across methods.
@@ V-C Ablation Study
Table III presents an ablation study of BiFlow on EgoTraj-TBD: Social Interaction (SI) within the contextual encoder, EgoAnchor (EA) distillation, and Shared Encoder (SE) used by both streams. We compare against MoFlow as a strong baseline, which can be viewed as a representative single-flow predictor that directly forecasts futures from past motion. Although it is not a strictly architecture-matched dual-versus-single control, BiFlow’s consistent improvements over MoFlow suggest that jointly modeling history denoising and future forecasting is beneficial under ego-view noise.
We then ablate components by enabling them progressively. When only SI is added, BiFlow achieves notable improvements over MoFlow, reducing minADE/minFDE by over 9% at KK=1. Incorporating EA further enhances performance, yielding a 13% improvement with the same $K$. With all components (SI, EA, and SE) integrated, the full model achieves the best performance, improving minADE and minFDE by 16%, 17%, and 16% at KK=1, KK=5, and KK=10, respectively. These results demonstrate that integrating social interaction cues and ego-centric intent patterns, particularly through EgoAnchor, significantly improves prediction accuracy and robustness under the deployment-realistic condition.
Figure 4: Visualization under ego-view noise. Gray: past trajectories (BEV and noisy FPV-derived histories). Orange: clean BEV futures (GT). Red: predicted futures.
@@ V-D Qualitative Results
Fig. 4 provides qualitative examples on EgoTraj-TBD. Although the BEV history and the FPV-derived history (both shown in gray) can differ substantially due to ego-view perception noise, BiFlow predicts future trajectories (red) that closely follow the clean BEV ground truth futures (orange). The examples highlight that BiFlow maintains accurate and physically plausible motion under deployment-realistic history corruption, indicating improved robustness to noisy and incomplete observations.
@@ VI Conclusion
While our approach shows competitive performance in the current setup, our benchmark relies on projecting FPV-derived tracks into metric space, which may introduce residual alignment error, and we do not yet quantify the impact of individual noise types. Generalization may also vary with camera geometry and sensing characteristics. Future work will improve calibration-robust projection, noise-stratified evaluation, and adaptation to diverse platforms and environments. We hope that EgoTraj-Bench and BiFlow will support the community with models more robust to the ego-centric observation challenges and progress toward reliable real-world deployment.